# Meta Pruning via Graph Metanetworks : A Meta Learning Framework for Network Pruning

### v1

@@ D.3 Full Results (Including Error Bars)
DCP [77] 76.01%(92.93%) 74.95%(-1.06%) 92.32%(-0.61%) 55.6%
CCP [49] 76.15%(92.87%) 75.21%(-0.94%) 92.42%(-0.45%) 54.1%
FPGM [19] 76.15%(92.87%) 74.83%(-1.32%) 92.32%(-0.55%) 53.5%
ABCP [42] 76.01%(92.96%) 73.86%(-2.15%) 91.69%(-1.27%) 54.3%
DMC [11] 76.15%(92.87%) 75.35%(-0.80%) 92.49%(-0.38%) 55.0%
Random [38] 76.15%(92.87%) 75.13%(-1.02%) 92.52%(-0.35%) 51.0%
DepGraph [9] 76.15%(-) 75.83%(-0.32%) - 51.7%
ATO [62] 76.13%(92.86%) 76.59%(+0.46%) 93.24%(+0.38%) 55.2%
DTP [39] 76.13%(-) 75.55%(-0.58%) - 56.7%
ours 76.14%(93.11%) 76.17%(+0.03%) 92.94%(-0.17%) 56.9%
ours 76.14%(93.11%) 76.27%(+0.13%) 93.05%(-0.06%) 56.9%
ours 76.14%(93.11%) 76.15%(+0.01%) 92.95%(-0.16%) 56.9%
@@ D.4 Computer Resources and Hyperparameters
Compute Resources GPU NVIDIA A100 40G
@@
Metanetwork num layer 4
hiddim 11
@@
node res ratio 0.001
edge res ratio 0.001
@@
Final Pruning metanetwork epoch 13
@@ Appendix E More About Ablation Study
E.1 Visualize "Test Acc VS. Speed Up" Curves
As mentioned in section 4.5, we visualize the "Test Acc VS. Speed Up" curves of metanetworks in Resnet56 on CIFAR10 in figure 4, VGG19 on CIFAR100 in figure 5 and ResNet 50 on ImageNet in figure 6.
Figure 4: Resnet56 on CIFAR10 Visualization
Figure 5: VGG19 on CIFAR100 Visualization
Figure 6: ResNet50 on ImageNet Visualization
E.2 Finetuning after Metanetwork
@@
Figure 7: Finetune After Metanetwork (ResNet56 on CIFAR10)
@@
$InitialPruning(Finetuning)\rightarrow Metanetwork(Finetuning)\rightarrow Pruning(Finetuning)$
@@
To illustrate this, we use the example of pruning ResNet56 on CIFAR10. In Figure 7(a), we visualize the "Test Accuracy vs. Speed Up" curves for models after the same metanetwork and after various amounts of finetuning. Without any finetuning after the metanetwork, the network’s accuracy is nearly zero (which is why there is no "finetune_0_epoch" line in the figure—it lies at the bottom). Directly proceeding to pruning and finetuning from such an unfinetuned network experimentally results in poor performance.
@@
  • Trend of Change: As the number of finetuning epochs increases, the flat portion of the curve becomes shorter, and the overall accuracy of this region improves.
@@
This suggests that the effectiveness of our method can be further improved by employing metanetworks trained for more meta-training epochs and applying additional finetuning afterward. In practice, however, we avoid excessive finetuning in order to keep the number of finetuning epochs within a reasonable range, allowing for multiple experimental trials.
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The Abstract and Introduction sections accurately reflect the contribution and scope of the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of our work is discussed in the conclusion section (5).
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We propose no theoretical result that needs to be proved.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We open sourced our code and provide everything you need to reproduce our code in Appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provides a URL link to our released code. All information about datasets and how to run our code can be found together with our code.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In Ablation Study and Appendix we talk about all the details of our experiments and how to choose hyparameter.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We run each experiment 3 times and calculate the error bars, which you can see in Appendix.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report our computer resources for each experiment in Appendix
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: This paper conforms to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: There is no societal impact of the work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper has no risks and needs no safeguards.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The assets utilized in this paper are appropriately acknowledged. The associated license and terms of use have been clearly stated and duly adhered to.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We open sourced our code and model with clear explanations.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We involve no crowdsourcing subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We involve no crowdsourcing subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We didn’t involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@ D.4.5 Hyperparameters
DCP (Zhuang et al., 2018) 76.01%(92.93%) 74.95%(-1.06%) 92.32%(-0.61%) 55.6%
CCP (Peng et al., 2019) 76.15%(92.87%) 75.21%(-0.94%) 92.42%(-0.45%) 54.1%
FPGM (He et al., 2019a) 76.15%(92.87%) 74.83%(-1.32%) 92.32%(-0.55%) 53.5%
ABCP (Lin et al., 2020b) 76.01%(92.96%) 73.86%(-2.15%) 91.69%(-1.27%) 54.3%
DMC (Gao et al., 2020) 76.15%(92.87%) 75.35%(-0.80%) 92.49%(-0.38%) 55.0%
Random (Li et al., 2022) 76.15%(92.87%) 75.13%(-1.02%) 92.52%(-0.35%) 51.0%
DepGraph (Fang et al., 2023) 76.15%(-) 75.83%(-0.32%) - 51.7%
ATO (Wu et al., 2024) 76.13%(92.86%) 76.59%(+0.46%) 93.24%(+0.38%) 55.2%
DTP (Li et al., 2023) 76.13%(-) 75.55%(-0.58%) - 56.7%
ours 76.14%(93.11%) 76.13%(-0.01%) 92.78%(-0.33%) 57.2%
ours 76.14%(93.11%) 76.24%(+0.20%) 93.09%(-0.02%) 57.1%
ours 76.14%(93.11%) 76.08%(-0.06%) 92.93%(-0.18%) 56.9%
@@
Compute Resources GPU NVIDIA A100 & NVIDIA RTX 4090
@@
Metanetwork num layer 6
hiddim 16
@@
node res ratio 0.002
edge res ratio 0.002
@@
Final Pruning metanetwork epoch 12
@@ Appendix E More About Ablation Study
E.1 Finetuning after Metanetwork
@@
Figure 9: Finetune After Metanetwork (ResNet56 on CIFAR10 as example)
@@
$\underbrace{\text{Initial Pruning (Finetuning)}}_{\text{Optional}}\rightarrow\underbrace{\text{Metanetwork (Finetuning)}\rightarrow\text{Pruning (Finetuning)}}_{\text{Necessary}}$ (29)
@@
To illustrate this, we use the example of pruning ResNet56 on CIFAR10. In Figure 9(a), we visualize the "Acc vs. Speed Up" curves for models after the same metanetwork and after various amounts of finetuning. Without any finetuning after the metanetwork, the network’s accuracy is nearly zero (which is why there is no "finetune_0_epoch" line in the figure—it lies at the bottom). Directly proceeding to pruning and finetuning from such an unfinetuned network experimentally results in poor performance.
@@
  • Trend of Change: As the number of finetuning epochs increases, the flat portion of the curve becomes shorter or stays the same, and the overall accuracy of this region improves.
@@
This suggests that the effectiveness of our method can be further improved by employing metanetworks trained for more meta-training epochs and applying more finetuning afterward. In practice, however, we avoid excessive finetuning in order to keep the number of finetuning epochs within a reasonable range, allowing for multiple experimental trials.
@@
Appendix F Experiments on Transferability
F.1 Transfer between datasets
The common pruning pipeline is:
$\text{Initial Pruning (Finetuning)}\rightarrow\text{Metanetwork (Finetuning)}\rightarrow\text{Pruning (Finetuning)}$ (30)
Pruning pipeline for None is:
Initial $Pruning (Finetuning)\text${Initial $\rightarrow\text{Pruning (Finetuning)}$ (31)
The dataset of a to be pruned network means we do all training and finetuning on it using this dataset. The dataset of a metanetwork means we generate dataset models, meta-train the metanetwork using this dataset.
We mainly have 2 hyperparameters here–initial speed up and final speed up . We approximately set them based on the difficulty of each dataset. Here dataset refers to the dataset of the to be pruned network (rows in Table 15). For CIFAR10, the initial pruning speed up is 1.32x and the final speed up is 3.0x. For CIFAR100 the initial pruning speed up is 1.32x and the final speed up is 2.5x. For SVHN, the initial pruning speed up is 3.0x and the final speed up is 10.0x.
Table 15: Transfer between datasets: All networks’ architecture is ResNet56. Columns represent the training datasets for the metanetwork, and rows represent the training datasets for the to be pruned network. “None” indicates using no metanetwork. Results with metanetwork is obviously better than no metanetwork (The only exception is when training datasets for the to be pruned network is SVHN, and we guess this is because the dataset SVHN itself is too easy).
[…]
CIFAR10 93.35 92.47 92.87 91.28
CIFAR100 69.97 70.16 69.25 68.91
SVHN 96.79 96.50 96.86 96.78
F.2 Transfer between Architectures
The common pruning pipeline is:
$\text{Initial Pruning (Finetuning)}\rightarrow\text{Metanetwork (Finetuning)}\rightarrow\text{Pruning (Finetuning)}$ (32)
Pruning pipeline for None is:
Initial $Pruning (Finetuning)\text${Initial $\rightarrow\text{Pruning (Finetuning)}$ (33)
The architecture of a to be pruned network means this network is constructed using this architecture. The architecture of a metanetwork means all dataset models we generated for meta-training this metanetwork use this architecture.
We mainly have 2 hyperparameters here–initial speed up and final speed up. We approximately set them based on the ability of each architecture. Here architecture refers to the architecture of the to be pruned network (rows in Table 16). For ResNet56, the initial pruning speed up is 1.32x and the final speed up is 3.0x. For ResNet110 the initial pruning speed up is 2.0x and the final speed up is 4.0x.
Table 16: Transfer between architectures. All training dataset is CIFAR10. Columns represent the architectures used for training the metanetwork, and rows represent the architecures of the to be pruned network. “None” indicates using no metanetwork. All results with metanetwork is obviously better than no metanetwork.
[…]
ResNet56 93.40 92.81 92.08
ResNet110 93.04 93.38 92.40
Appendix G Expriments on Transformers
G.1 MHSA to Graph
[…]
$Q_{h}=\mathbf{X}\mathbf{W}_{h}^{Q},\quad K_{h}=\mathbf{X}\mathbf{W}_{h}^{K},\quad V_{h}=\mathbf{X}\mathbf{W}_{h}^{V}.$
Then each head computes attention via
$Y_{h}=\text{softmax}(Q_{h}K_{h}^{\top})V_{h},$
The outputs $Y_{1},\ldots,Y_{H}$ are concatenated and linearly projected:
$\text{MHSA}(\mathbf{X})=\text{Concat}(Y_{1},\ldots,Y_{H})\mathbf{W}^{O}.$
The input and output are both $d$-dimensional. Each head produces $d_{H}$-dimensional outputs. When changing to a graph, we represent this with $d$ input nodes in the first layer, $H\cdot d_{H}$ attention head nodes in the second layer, and $d$ output nodes in the third layer.
Between the first and the second layer, we model the three projection types (query, key, value) using multidimensional edge features: for each edge $(i,j)$ in head $h$, the feature is
$e_{ij}^{h}=\big((\mathbf{W}_{h}^{Q})_{ij},\;(\mathbf{W}_{h}^{K})_{ij},\;(\mathbf{W}_{h}^{V})_{ij}\big).$
Between the second and third layer, concatenation and the final projection $\mathbf{W}^{O}$ are handled naturally by the graph structure and treated as a standard linear layer.
In summary, every pair of nodes between the first and second layers is connected by an edge characterized by three-dimensional features, which represent the attention parameters. Each pair of nodes between the second and third layers is connected by an edge with a single-dimensional feature, corresponding to the final linear projection.
G.2 Equivalent Conversion between Network and Graph
We change ViT-B/16 into a graph with node featrues of 6 dimensions and edge features of 1, 3 or 256 dimensions. We employ 3 linear layers to project edge features of these 3 different dimensions into the same hidden dimension, enabling their integration into our graph neural network.
The node features consist of 6 features, they are weight and bias of previous layer norm, bias of previous linear layer, biases of query, key, value of previous attention layer. In the case that any of these features doesn’t exist, they are replaced by a default value of $[1,0,0,0,0,0]$ at their corresponding positions.
The edge features are constructed almost in the same way as previous experiments. The only new part is MHSA, and we construct it as described in Appendix G.1.
G.3 Meta Training
Data Models: We use the default ViT-B/16 provided by PyTorch as the only data model. All training and pruning are conducted on this model, as training several separate models from scratch would be prohibitively time-consuming. Moreover, this choice ensures consistency with prior work in the field. Based on the Acc VS. speed up curve of the original ViT (Figure 10), we do initial pruning with a speed up of 1.0370x and apply no finetuning. This configuration yields a network with an accuracy of 81%.
Figure 10: Acc VS. Speed Up curve of the origin ViT
One Meta Training Epoch : The same as ResNet50 on ImageNet (Appendix D.4.2).
Training: We train our metanetwork with learning rate 0.01, milestone "3", weight decay 0.0005 and pruner reg 1000. Finally we use metanetwork from epoch 6 as our final network for pruning.
G.4 GPU Usage
Meta-Training and feed forward through the metanetwork during pruning requires large VRAM and needs to be run on NVIDIA A100. All other training and finetuning can be run on NVIDIA RTX 4090. In practice, we use 8 gpus in parallel.
G.5 Full Results
All results are summarized in Table 17. Note that unlike CNNs, Vision Transformers naturally lacks inductive bias and need much more epochs of finetuning to reach higher accuracy. For other works like DepGraph (Fang et al., 2023), theirs method need to finetune 600 epochs in total, among them, 300 epochs are sparsity learning (slower than usual finetuning) and it takes more than a week to run a single round. We don’t have enough time to fully run the experiments with more finetuning for now, but we will continue tune our hyperparameters, add more epochs and hopefully give a better results in the next version of our paper.
G.6 Hyperparameters
See Table 18.
Table 17: ViT-B/16 on ImageNet
[…]
ViT-B/16 (Dosovitskiy et al., 2021) 81.07% - - 17.6
CP-ViT (Song et al., 2022) 77.91% 77.36% -0.55% 11.7
DepGraph (Fang et al., 2023) 81.07% 79.17% -1.90% 10.4
MetaPruning(ours) 81.07% 77.59% -2.83% 11.0
Table 18: ViT-B/16 on ImageNet Hyperparameters
Types Name Value
Compute Resources GPU NVIDIA A100 & NVIDIA RTX 4090
parallel 8
Batch size small batch size 128 $\times$ 8
big batch size 32 $\times$ 8 $\times$ 200
Prepare Data Models data model num 1
Generate Data Model epoch 0
lr -
weight decay -
milestone -
Initial Pruning speed up 1.0370
Finetuning epoch 0
lr -
weight decay -
milestone -
Metanetwork num layer 3
hiddim 4
in node dim 6
node res ratio 0.1
edge res ratio 0.1
Meta Training lr 0.01
weight decay 0.0005
milestone "3"
pruner reg 1000
Final Pruning metanetwork epoch 6
speed up 1.6000
Finetuning After Metanetwork epoch 100
lr 0.0001
weight decay 0.01
scheduler cosineannealinglr
Finetuning After Pruning epoch 100
lr 0.0001
weight decay 0.0001
scheduler cosineannealinglr
Appendix H The Use of Large Language Models
We only use LLMs to check grammar errors and polish our writing. All other works are done by human writers.